OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3 x = –5 and x = –11 x = 5 and x = 11 x = –5 and x = 11 x = 5 and x = –11 help me please

OpenStudy (cwrw238):

|–2x + 16| – 9 = –3 |–2x + 16| = 6 means -2x + 16 = 6 also -2x + 16 = -6 solve these 2 equations for x and your are home